OpenStudy (xoxolove456):

Madison has a garden in her back yard that is 7 feel longer that it is wide. She also has a path around the entire garden that is 3 feet wide. The area of the garden and the path around the garden is 450 square feet. What is the length of the garden?

OpenStudy (welshfella):

let length of the garden = x feet then the width = x - 7 feet

OpenStudy (welshfella):

Now we need to find an expression for the area of the garden + the path length = x + 2(3) and width = (x - 7) + 2(3) = x + 6 and x - 1

OpenStudy (welshfella):

so we can set up an equation for the total area of garden + path:- (x + 6)(x - 1) = 450 solve for x

OpenStudy (welshfella):

ignore any negative roots

OpenStudy (xoxolove456):

So I just solve the last one?